OpenStudy (anonymous):

Write the missing numbers. 1/6 = 7/? 3/4 = 15/? 11/12 = 44/? 4/7 = 12/? 2/3 = ?/? 3/4 = ?/? 9/10 = ?/? 5/9 = ?/?

OpenStudy (anonymous):

hint... \[\LARGE \frac{1}{6}=\frac{1\cdot 7}{6\cdot 7}= \frac{7}{42}\] \[\LARGE \frac{3}{4}=\frac{3\cdot 5}{4\cdot 5 }= \frac{15 }{?}\] keep going...

OpenStudy (anonymous):

well I don't really know what kind of rule is going on... since last four fractions are really confusing me O_O !!

OpenStudy (anonymous):

It says to find the equivalent fraction for the last four. : )

OpenStudy (anonymous):

ahhh... so it means you can type every fraction that you want. But it has to be equivalent.. ahhah I get it now.. multiply or simplify(if you can) numerator and denominator. Both together. if you multiply numerator by a number you should multiply denominator as well with the same number. ;) ... do you think you can do it ?

OpenStudy (anonymous):

Hint.. \[\LARGE \frac{11}{12}=\frac{11\cdot 4}{12\cdot 4}=\frac{44}{48}\] \[\LARGE \frac{4}{7}=\frac{4\cdot 3}{7\cdot 3}=\frac{12}{?}\] \[\LARGE \frac{2}{3}=\frac{2\cdot 2}{3\cdot 2}=\frac{?}{?}\] keep going ..

OpenStudy (anonymous):

@MusicalAnimal Still waiting? ... do you understand them ?

OpenStudy (anonymous):

Yes. Thank you so much!

OpenStudy (anonymous):

if you don't understand them ...tell me. don't feel bad, just let me know, I'll do them all !